OpenStudy (anonymous):

Find the vertex, the line of symmetry the maximum or minimum value of the quadratic function, and graph the function. f(x)=-2x^2+2x+9

OpenStudy (anonymous):

I got the answer (1/2 , 19/2) and I need to simplify so I put (1/2, 8) and that was not it

OpenStudy (anonymous):

This is Algebra...I believe

OpenStudy (anonymous):

this involves differentiation.. ! to find maximas and minimas

OpenStudy (anonymous):

yes, I am taking math 117 at phoenix

OpenStudy (anonymous):

i bet jimmy is giving a huge lecture..so just wait for his reply :D

OpenStudy (anonymous):

vertex is -2 / -4 = 1/2,y y = -2*1/4 + 1 + 9 = 19/2 vertex = (1/2 , 19/2) or (0.5,9.5) line of symmetry is x = 1/2 term in x^2 is negative so function has a maximum this is at (1/2, 19/2) or (0.5,9.5)

OpenStudy (anonymous):

To get maxima or minima differentiatie so you ll get\[-4x+2=0\] this gives x = (1/2).. substitute this value and you ll get your maxima or minima

OpenStudy (anonymous):

maximum value of the function is 9.5